Department of Electrical Engineering, BUKC have purchased MatLab software for Rs. 1,000,000 now an annual payments of Rs. 100,000 per year for 6 years starting 3 years from now for annual upgrades. What is the present worth of the payments if the interest rate is 7% per year?
1
Expert's answer
2020-05-19T09:58:42-0400

The present worth of the payments if the interest rate is 7% per year is:

"PW = 1,000,000 + \frac{100,000}{1.07^3} + \frac{100,000}{1.07^4} + \frac{100,000}{1.07^5} + \frac{100,000}{1.07^6} + \frac{100,000}{1.07^7} + \frac{100,000}{1.07^8} = 1,416,328.03."